# Undersmoothed LASSO Models for Propensity Score Weighting and Synthetic Negative Control Exposures for Bias Detection

### v2

@@ 4 Results
Figures 2 and 3 further show that general patterns in bias between the synthetic negative control studies and study population were similar (both in terms of bias, shown in Plots A and B, and relative or percent bias, shown in Plots C and D). However, while overall patterns in bias were similar, the magnitude of bias within the synthetic negative control studies (synthetic bias) was smaller than the bias in the study population (Plots A and B). As the sample size increased, differences in the magnitude of bias between the synthetic cohorts vs the actual study population decreased.
@@
Figure 6: Absolute standardized difference of each covariate across exposure groups for Simulation Setup 1. Plots A through D correspond to sample sizes 5K (Plot A), 10K (Plot B), 20K (Plot C), and 40K (Plot D). The x-axis for each plot shows standardized differences before PS weighting (unadjusted), while the y-axis shows the standardized differences after PS weighting. The red horizontal dotted line indicates the value that is commonly used to determine adequate covariate balance after PS weighting (standardized difference $<0.1$). For the plots shown here, weighting was done using inverse probability weights and the Lasso model used to estimate the PS was tuned using cross validation.
@@ 5 Discussion
In this study, we considered using balance criteria to determine the degree of undersmoothing when fitting LASSO models for PS weighting. Because no tuning criteria and weighting approach are univerally best, we further proposed a framework to generate synthetic negative control exposures to detect bias caused by improper undersmoothing for the given study and PS weighting approach. Numerical experiments suggest that using balance criteria to undersmooth LASSO models can reduce bias in PS weighted estimators compared to estimators that are tuned using cross-validation. Numerical studies further suggest that synthetic negative control exposures can be useful for bias detection.
@@
Finally, it is important to highlight that future work could explore variations of the proposed framework for generating synthetic negative control exposures. In particular, the proposed algorithm can be flexible in terms of what sampling technique is applied. Here, we proposed the use of bootstrap oversampling to generate samples of the same size as the original study cohort on which to evaluate estimators. However, if one is concerned about evaluating estimators in samples that contain many repeated observations, future work could consider alternative subsampling techniques, such as the $m$-out-of-$n$ bootstrap or $m$-out-of-$n$ subsampling without replacement [51, 52]. Subsampling techniques are widely regarded as being more robust than the traditional $n$-out-of-$n$ bootstrap (and bootstrap oversampling) as they are asymptotically valid under weaker conditions [51, 52, 32]. But subsampling techniques require one to select the size of the subsamples which involves some exercise of judgment and are limited in that the size of the subsamples must be smaller than the original study cohort. Future work could also consider variations that re-estimate the CV LASSO used to generate synthetic exposure probabilities within each bootstrapped sample.
@@ Acknowledgements
This work was funded by National Institutes of Health grant NIH RO1LM013204 and Patient-Centered Outcomes Research Institute contract PCORI ME-2022C1-25646.
@@
HAL is a machine learning prediction algorithm that can be used to nonparametrically estimate regression functions. To understand HAL, it is necessary to understand the conditions (or assumption) required by HAL. First, HAL requires the total variation of a function (or variation norm) to be bounded, where the total variation can be thought of as a measure of the complexity of a function. For example, in the simple case of a monotone function, $f(\cdot)$, on the interval $[0,1]$, the total variation is simply $|f(1)-f(0)|$. For more general functions, the total variation is equal to the cumulative sum of all the absolute incremental changes in the function’s values over its domain. Intuitively, if we plotted a non-linear/non-monotone continuous univariate function and traced the function with a piece of string, the total variation of the function can be thought of as the length of the string. Being able to bound the total variation of a function ensures that the function cannot wiggle around too much. The concept of bounded total variation can be extended to higher dimensions (and non-continuous functions), but the general idea remains the same.
@@
$logit(g(W))=\gamma_{0}+W^{\top}\gamma$ (4)
@@
Proposition 1: Assume we have the data structure, $\{Y,X,A,Z\}$, where $Z$ is a negative control exposure. If $Y^{a}\perp\!\!\!\perp A|X$ it follows that $Y\perp\!\!\!\perp Z|X$.
@@
Let $Y^{a}\perp\!\!\!\perp A|X$. From the definition of a negative control exposure, $P(Z|X)$ can be written as a 1-to-1 function of $P(A|X)$, and there is no unmeasured common cause of $Z$ and $Y$. This implies that $X$ is sufficient to satisfy $Y^{z}\perp\!\!\!\perp Z|X$. From the definition of a negative control exposure, we further know that $Y^{(a,z)}=Y^{a}$ (no causal effect of $Z$ on $Y$). This further implies that $Y^{z=0}=Y^{z=1}=Y$. Therefore, $Y\perp\!\!\!\perp Z|X$
@@
Proposition 2: Assume we have the data structure, $\{Y,X,A,Z\}$, where $Z$ is a negative control exposure. If $Y^{a=0}\perp\!\!\!\perp A|X$ it follows that $Y\perp\!\!\!\perp Z|X,A=0$.
@@
Let $Y^{a=0}\perp\!\!\!\perp A|X$. From the definition of a negative control exposure, $Y^{a}\perp\!\!\!\perp Z|X$ and $A\perp\!\!\!\perp Z|X$. This implies that $Y^{a=0}\perp\!\!\!\perp Z|X,A=0$. Therefore, $Y\perp\!\!\!\perp Z|X,A=0$ by consistency. ∎
S1.4   Proof for Comment 2 on Definition 2 in Section 2.4.1
@@
If we make the simplifying assumption that the relationship between $X$ and $A$ is linear (or linear on the log-odds scale), the this implies that $P(X|A=1)-P(X|A=0)$ does not vary between strata of $X$. Under this condition, Vanderweele & Arah (2011) show that the above expression then simpifies to

### v3

@@ 4 Results
Figures 2 and 3 further show that general patterns in bias between the synthetic negative control studies and study population were similar (both in terms of bias, shown in Plots A and B, and relative or percent bias, shown in Plots C and D). However, while overall patterns in bias were similar, the magnitude of bias within the synthetic negative control studies (synthetic bias) was smaller than the bias in the study population (Plots A and B). As the sample size increased, differences in the magnitude of bias between the synthetic cohorts vs the actual study population decreased. Supplemental Figures 1 and 2 further show that for both simulation setups, the undersmoothed LASSO PS weighted estimators resulted in coverage rates that were closer to the nominal confidence level compared to the PS weighted estimators that were tuned using cross-validation.
@@
Figure 6: Absolute standardized difference of each covariate across exposure groups for Simulation Setup 1. Plots A through D correspond to sample sizes 5K (Plot A), 10K (Plot B), 20K (Plot C), and 40K (Plot D). The x-axis for each plot shows standardized differences before PS weighting (unadjusted), while the y-axis shows the standardized differences after PS weighting. The red horizontal dotted line indicates the value that is commonly used to determine adequate covariate balance after PS weighting (standardized difference <0.1<0.1). For the plots shown here, weighting was done using inverse probability weights and the LASSO model used to estimate the PS was tuned using cross validation.
@@ 5 Discussion
In this study, we considered using balance criteria to determine the degree of undersmoothing when fitting LASSO models for PS weighting. Because no balance criterion for tuning LASSO PS weighted estimators is universally best, we further proposed a framework to generate synthetic negative control exposures to detect bias caused by improper undersmoothing. Numerical experiments suggest that using balance criteria to undersmooth LASSO models can reduce bias in PS weighted estimators compared to estimators that are tuned using cross-validation. Numerical studies further suggest that synthetic negative control exposures can be useful for bias detection.
@@
Finally, it is important to highlight that future work could explore variations of the proposed framework for generating synthetic negative control exposures. In particular, the proposed algorithm can be flexible in terms of what sampling technique is applied. Here, we proposed the use of bootstrap oversampling to generate samples of the same size as the original study cohort on which to evaluate estimators. However, if one is concerned about evaluating estimators in samples that contain many repeated observations, future work could consider alternative subsampling techniques, such as the $m$-out-of-$n$ bootstrap or $m$-out-of-$n$ subsampling without replacement [53, 54]. Subsampling techniques are widely regarded as being more robust than the traditional $n$-out-of-$n$ bootstrap (and bootstrap oversampling) as they are asymptotically valid under weaker conditions [53, 54, 36]. But subsampling techniques require one to select the size of the subsamples which involves some exercise of judgment and are limited in that the size of the subsamples must be smaller than the original study cohort. Future work could also consider variations for fitting the cross-validated LASSO that is used to generate synthetic exposure probabilities. Alternative possibilities that merit future consideration could include using a cross-validated relaxed LASSO [62], as well as repeatedly sampling coefficient vectors from posterior distributions to put less weight on the quality of a specific estimate for the PS (in practice this might be accomplished by fitting the cross-validated LASSO to bootstrapped samples).
@@ Acknowledgements
This work was funded by National Institutes of Health grant NIH R01LM013204 and Patient-Centered Outcomes Research Institute contract PCORI ME-2022C1-25646.
@@
The Highly Adaptive Lasso (HAL) is a machine learning prediction algorithm that can be used to nonparametrically estimate regression functions. To understand HAL, it is necessary to understand the conditions (or assumption) required by HAL. First, HAL requires the total variation of a function (or variation norm) to be bounded, where the total variation can be thought of as a measure of the complexity of a function. For example, in the simple case of a monotone function, $f(\cdot)$, on the interval $[0,1]$, the total variation is simply |f(1)f(0)||f(1)-f(0)|. For more general functions, the total variation is equal to the cumulative sum of all the absolute incremental changes in the function’s values over its domain. Intuitively, if we plotted a non-linear/non-monotone continuous univariate function and traced the function with a piece of string, the total variation of the function can be thought of as the length of the string. Being able to bound the total variation of a function ensures that the function cannot wiggle around too much. The concept of bounded total variation can be extended to higher dimensions (and non-continuous functions), but the general idea remains the same.
@@
$logit(g(W))=\gamma_{0}+W^{\top}\gamma,$ (4)
@@
Proposition 1: Assume we have the data structure, $\{Y,X,A,Z\}$, where $Z$ is a negative control exposure that satisfies Definition 1. For any $X_{s}\subseteq X$, if $(Y^{a=1},Y^{a=0})\perp\!\!\!\perp A|X_{s}$ then $Y\perp\!\!\!\perp Z|X_{s},A$.
@@
Let $X_{s}$ be any subset of $X$ such that $Y^{a}\perp\!\!\!\perp A|X_{s}$. The proportional odds assumption in Defininition 1 ensures that $P(Z|X)$ can be written as a 1-to-1 function of $P(A|X)$. This implies that all predictors of $Z$ and $A$ are equivalent. Therefore, for any variable not in $X_{s}$ that is a common cause of $Z$ and $Y$, that variable’s effect on $Y$ must be mediated through the exposure, $A$. Otherwise, $X_{s}$ would not be sufficient to satisfy $Y^{a}\perp\!\!\!\perp A|X_{s}$. This implies that $X_{s}$ is sufficient to satisfy $Y^{z}\perp\!\!\!\perp Z|X_{s},A$. From the definition of a negative control exposure, we further know that $Y^{(a,z)}=Y^{a}$ (no causal effect of $Z$ on $Y$). This further implies that $Y^{z=0}=Y^{z=1}=Y$. Therefore, $Y\perp\!\!\!\perp Z|X_{s},A$. ∎
@@
Proposition 2: Assume we have the data structure, $\{Y,X,A,Z\}$, where $Z$ is a negative control exposure that satisfies Definition 1. For any $X_{s}\subseteq X$, if $Y^{a=0}\perp\!\!\!\perp A|X_{s}$ then $Y\perp\!\!\!\perp Z|X_{s},A=0$.
@@
Let $X_{s}$ be any subset of $X$ such that $Y^{a=0}\perp\!\!\!\perp A|X_{s}$. From the definition of a negative control exposure, $Y^{a}\perp\!\!\!\perp Z|X_{s},A$ and $A\perp\!\!\!\perp Z|X_{s}$. This implies that $Y^{a=0}\perp\!\!\!\perp Z|X_{s},A=0$. Therefore, $Y\perp\!\!\!\perp Z|X,A=0$ by consistency. ∎
S1.4   Further Discussion of the Proportional Odds Assumption
The goal of the proportional odds assumption is to have covariate differences across negative control exposure groups within the unexposed that approximate covariate differences across the exposure groups in the full population. If covariate differences across exposure groups match those across negative control exposure groups within the unexposed, then under the assumption that the association between $X$ and $A$ is linear, it can be shown that the magnitude of confounding bias caused by $X$ on the relationship between $A$ and $Y^{a=0}$, will be equal to the bias caused by $X$ on the relationship between $Z$ and the observed outcome, $Y$, among the unexposed on the absolute scale (see formal description of this statement and proof below). If the relationship between $X$ and $A$ is not linear, but is linear in some expanded basis (e.g., the Highly Adaptive Lasso), then the same arguments hold and are generally applicable by replacing $X$ with the expanded set of basis functions. In practice, however, the proportional odds assumption only guarantees that covariate differences across negative control exposure groups within the unexposed will approximate differences across exposure groups in the full study population; they will not be exact.
@@
If we make the simplifying assumption that the relationship between $X$ and $A$ is linear (or linear on the log-odds scale), the this implies that P(X|A=1)P(X|A=0)P(X|A=1)-P(X|A=0) does not vary between strata of $X$. Under this condition, Vanderweele & Arah (2011) show that the above expression then simpifies to